EXCESS CAPACITY Williamson Industries has $5 bilion in sales and $1.7 billion in fixed assets. Currently, the companys fixed assets are operating at 90% of capacity. a. What level of sales could Williamson Industries have obtained if it had been operating at full capacity? Write out your answer completely. For example, 25 billion should be entered as 25,000,000,000. Round your answer to the nearest cent. 5560000000 b. What is Williamsons target fixed assets/sales ratio? Round your answer to two decimal places. 30.63 C. If Williamsons sales increase 12%, how large of an increase in fixed assets will the company need to meet its target fixed assets/sales ratio? Write out your answer completely. For example, 25 billion should be entered as 25,000,000,000. Round your answer to the nearest cent. Negative amount should be indicated by a minus sign. Do not round intermediate calculations. 15280000

pleaseeee help with C only, i have tried everything!!:(
Answer #1

c.

new sales=(5*1000000000)*(1+12%)=5600000000

no increase in fixed assets up to sales of =5000000000/90%=5555555555.56

increase in fixed assets=30.6%*(5600000000-5555555555.56)

=13600000

the above is answer..